# Transfer Learning for Spatial Autoregressive Models

### v1

@@ Abstract
The spatial autoregressive (SAR) model has been widely applied in various empirical economic studies to characterize the spatial dependence among subjects. However, the precision of estimating the SAR model diminishes when the sample size of the target data is limited. In this paper, we propose a new transfer learning framework for the SAR model to borrow the information from similar source data to improve both estimation and prediction. When the informative source data sets are known, we introduce a two-stage algorithm, including a transferring stage and a debiasing stage, to estimate the unknown parameters and also establish the theoretical convergence rates for the resulting estimators. If we do not know which sources to transfer, a transferable source detection algorithm is proposed to detect informative sources data based on spatial residual bootstrap to retain the necessary spatial dependence. Its detection consistency is also derived. Simulation studies demonstrate that using informative source data, our transfer learning algorithm significantly enhances the performance of the classical two-stage least squares estimator. In the empirical application, we apply our method to the election prediction in swing states in the 2020 U.S. presidential election, utilizing polling data from the 2016 U.S. presidential election along with other demographic and geographical data. The empirical results show that our method outperforms traditional estimation methods.
keywords:
Spatial autoregressive model, transfer learning, U.S. presidential election
journal: Journal of Econometrics
@@ 1 Introduction
Spatial dependence among spatial units exists widely in various fields of economics, including urban, real estate, public, agricultural, environmental economics, finance, and so on. The spatial autoregressive (SAR) model (Cliff and Ord,, 1973; Ord,, 1975; Cliff and Ord,, 1975) and its variants are extensively applied in numerous empirical studies concerning spatial competition and spatial spillover effects. For instance, the SAR model is employed to investigate the crime rates in 49 areas of Columbus, Ohio (Anselin,, 1988), budget spillover effects and policy interdependence (Case et al.,, 1993), cigarette demands in various states in the U.S. (Baltagi and Li,, 2004), agricultural land prices in Northern Ireland (Kostov,, 2009), the mutual influence on the stock returns of companies with the same headquarters location (Pirinsky and Wang,, 2006) and the group effect of the dividend policy of listed companies (Grennan,, 2019). Besides, there are also large amounts of applications in other fields where spatial correlations are considered, including political studies (Peters et al.,, 2004), ecology (Ver Hoef et al.,, 2018) and climate sciences (Okunlola et al.,, 2021).
@@
When the sample size of the target data is limited, it is possible to obtain less precise estimates for most statistical models. This issue is particularly relevant in spatial econometrics, where sample sizes can often be quite limited. For instance, in the study investigating crime rates in Columbus, Ohio, the sample size is restricted to just 48 (Anselin,, 1988). This limitation underscores the necessity for additional research aimed at leveraging information from similar datasets to elevate estimation performance. Consequently, there is a pressing need for the exploration of methods to improve the estimation performance of the SAR models through the incorporation of information derived from similar source datasets. Transfer learning is a useful approach to incorporate information from source datasets (Pan and Yang,, 2009; Torrey and Shavlik,, 2010; Zhuang et al.,, 2021; Hu and Zhang,, 2023). The main idea of transfer learning is that the knowledge acquired in one context can be transferred to improve the performance of estimation or prediction when applied to another, even if the two tasks or domains are not identical. In transfer learning, there are typically two main components: (1) The first one is the target domain, to which the knowledge is transferred. The target domain is of our interest but particularly its sample size is limited. (2) The second one is the source domain, from which the knowledge is transferred. In general, the model is pre-trained using the source datasets.
@@
In the recent literature, Bastani, (2021) introduced a two-step estimator that employs high-dimensional techniques to capture biases between the target and source data and has provided an upper bound for the proposed estimators. Li et al., 2022b explored more general scenarios under weaker assumptions and proposed a data-driven method for detecting similarity between the sources and the target. Tian and Feng, (2023) extended the linear model framework to the generalized linear model and introduced an entropy-based method for identifying transferable sources. For more reference on related works, see Cai and Pu, (2019), Cai et al., (2022, 2023), and Li et al., 2022c ; Li et al., 2022a .
@@
In this paper, we introduce a transfer learning framework under the SAR model (tranSAR for short) to enhance the performance of both estimation and prediction. Our study makes three contributions from methodology, statistical theory and empirical application in the literature of the SAR models. First, in methodology, we introduce a two-stage algorithm ($\mathcal{A}$-TranSAR), consisting of a transferring stage and a debiasing stage, to estimate the unknown parameters when the informative source data indexed by $\mathcal{A}$ are known. If we do not know which sources to transfer, a transferable source detection algorithm is proposed to detect informative sources data based on spatial residual bootstrap to retain the necessary spatial dependence. To the best of our knowledge, this is the first work to introduce the transfer learning to the SAR models. Second, in statistical theory, we establish the theoretical convergence rates for the $\mathcal{A}$-TranSAR estimators under the SAR model and theoretically show that the transfer learning can improve the estimation accuracy with the help of the information from the transferable source data. We also derive the detection consistency of the transferable source detection algorithm to avoid the negative transfer when which sources to transfer is unknown. Third, in the empirical application, we apply our method to predict the election outcomes in swing states in the 2020 U.S. presidential election via utilizing polling data from the 2016 U.S. presidential election along with other demographic and geographical data. The empirical results show that our method outperforms traditional spatial econometrics methods.
In the rest of our paper, Section 2 introduces a transfer learning framework for the SAR model and its estimation when the informative source sets are known. When the transferable source data are unknown, a source detection algorithm is also proposed based on spatial residual bootstrap in this section. Section 3 establishes the estimation convergence rates of the transfer learning estimators for the SAR model and derive the detection consistency of the transferable source detection algorithm. In Section 4, we investigate the numerical performance of the proposed methods using Monte Carlo simulations. Section 5 applies the proposed algorithms to the empirical application on U.S. presidential election prediction. We conclude the paper in Section 6.
@@
2 Transfer Learning for Spatial Autoregressive Model
2.1 Model setting
We consider the spatial autoregressive (SAR) model for the target data,
$Y^{(0)}=\sum_{l=1}^{p}\lambda_{l0}W_{l}^{(0)}Y^{(0)}+\sum_{j=1}^{q}X^{(0)}_{j}\beta_{j0}+V^{(0)},$ (2.1)
@@ 2.1 Model setting
where $Y^{(0)}$ is an $n_{0}\times 1$ vector of spatially correlated dependent variables, $W_{l}^{(0)}$ is an $n_{0}\times n_{0}$ non-stochastic spatial weight matrix for each $l\in[p]$, $X_{j}^{(0)}$ is an $n_{0}\times 1$ vector of exogenous regressors independent of the error term $V^{(0)}$ for each $j\in[q]$, $V^{(0)}$ is an $n_{0}\times 1$ vector of independently and identically distributed (i.i.d.) random variables with mean 00 and variance $\sigma^{2}$. Let $\lambda_{0}={(\lambda_{10},\dots,\lambda_{p0})}^{\intercal}$, $\beta_{0}=(\beta_{10},\dots,\beta_{q0})^{\intercal}$ and $\theta_{0}\equiv$ ${({\lambda_{0}}^{\intercal},{\beta_{0}}^{\intercal})}^{\intercal}$.
@@
Similar to general transfer learning settings as in Bastani, (2021)Li et al., (2021)Tian and Feng, (2023), etc., in addition to the target study (2.1), additional samples are obtained from $K$ auxiliary studies (also termed as source datasets), where $K$ is assumed to be fixed. For the $k$-th source dataset, we also consider the SAR model,
$Y^{(k)}=\sum_{l=1}^{p}\lambda_{l0}^{(k)}W_{l}^{(k)}Y^{(k)}+\sum_{j=1}^{q}X^{(k)}_{j}\beta^{(k)}_{j0}+V^{(k)},\ k\in[K],$ (2.2)
@@
$\mathcal{A}\equiv\mathcal{A}_{h}=\{0\leq k\leq K:\|\delta^{(k)}\|_{1}\leq h\},$ (2.3)
@@ 2.2 Estimation with known informative source data
$\widehat{\omega}=\underset{\omega\in\mathbb{R}^{p+q}}{\arg\min}\left\{\frac{1}{2n_{\mathcal{A}}}\sum_{k\in\mathcal{A}}\operatorname{\mathit{\ell}}_{1}\left(\mathfrak{D}^{(k)}\mid\omega\right)+\lambda_{\omega}\|\omega\|_{1}\right\},$ (2.4)
where $\operatorname{\mathit{\ell}}_{1}(\cdot\mid\omega)$ is a generic loss function with parameter $\omega$, and $n_{\mathcal{A}}\equiv\sum_{k\in\mathcal{A}}n_{k}$ is the total sample size of all informative source data. For the SAR model, the estimator could be the two-stage least squares (2SLS) estimator (Kelejian and Prucha,, 1998; Lee,, 2003), generalized method of moments estimator (GMM) (Kelejian and Prucha,, 1999; Lee,, 2007; Liu and Lee,, 2010), quasi-maximum likelihood estimator (Lee,, 2004, QMLE) and so on. The main requirement for the first step estimator is a suitable convergence rate discussed in detail in Section 3. Here the tuning parameter $\lambda_{\omega}$ could be zero in low dimensional cases. When the dimension of parameters is greater than the sample size in high dimensional cases, the penalized SAR estimation with nonzero $\lambda_{\omega}$ could produce a sparse estimator, (for example, Higgins and Martellosio, (2023)). See also Zhang and Yu, (2018) for spatial weight matrices selection with model averaging.
@@
The index set $\mathcal{A}$ may or may not contain the index 00. This depends on the actual application scenarios. In some cases the source data are available, and in some cases, it may be difficult to obtain due to data privacy and security considerations. In the latter case, the owner of the source data can provide the estimation results obtained by the first stage learner without the target data instead of the original source data, and the target task will use this result for the next debiasing stage.
@@
$\underset{\delta\in\mathbb{R}^{p+q}}{\arg\min}\frac{1}{2n_{0}}\operatorname{\mathit{\ell}}_{2}\left(\mathfrak{D}^{(0)}\mid\widehat{\omega}+\delta\right)+\lambda_{\delta}\|\delta\|_{1}.$ (2.5)
@@
$\widehat{\theta}_{\mathcal{A}\textit{-TranSAR}}=\widehat{\omega}+\widehat{\delta}.$ (2.6)
@@
Input: target dataset: $\mathfrak{D}^{(0)}$; sources datasets: $(\mathfrak{D}^{(k)},k\in\mathcal{A})$; transferring set: $\mathcal{A}$; and tuning parameters: $\lambda_{\omega}$, and $\lambda_{\delta}$

### v2

# tranSAR_supp

@@ Abstract
It is important to incorporate spatial geographic information into U.S. presidential election analysis, especially for swing states. The state-level analysis also faces significant challenges of limited spatial data availability. To address the challenges of spatial dependence and small sample sizes in predicting U.S. presidential election results using spatially dependent data, we propose a novel transfer learning framework within the SAR model, called as tranSAR. Classical SAR model estimation often loses accuracy with small target data samples. Our framework enhances estimation and prediction by leveraging information from similar source data. We introduce a two-stage algorithm, consisting of a transferring stage and a debiasing stage, to estimate parameters and establish theoretical convergence rates for the estimators. Additionally, if the informative source data are unknown, we propose a transferable source detection algorithm using spatial residual bootstrap to maintain spatial dependence and derive its detection consistency. Simulation studies show our algorithm substantially improves the classical two-stage least squares estimator. We demonstrate our method’s effectiveness in predicting outcomes in U.S. presidential swing states, where it outperforms traditional methods. In addition, our tranSAR model predicts that the Democratic party will win the 2024 U.S. presidential election.
Keywords: Spatial autoregressive model, transfer learning, U.S. presidential election, spatial data analysis, high dimensional analysis.
@@ 1 Introduction
In recent years, amounts of Geographic Information Science (GIS) data have been used in election campaigns (Peters et al.,, 2004). Geography is the common denominator of where voters live, how residents engage in the voting process, how officials manage elections, and where elections campaign runs. Based on population data and geographic information, election district boundaries are resettled to draft multiple proposed plans and demonstrate the effects of several approaches. Political campaigns make full use of geographic information and related tools to obtain underlying support from voters, especially the ones from swing states, which are pivotal in determining the outcome of elections. Swing states, also known as battleground states, are crucial because they do not consistently support one party, making their election outcomes unpredictable and highly influential in election results. Shaping voter preference in swing states is an essential problem, especially for prediction. Including spatial information in data analysis brings a novel insight into the prediction. There are at least two challenges in using spatial information for predicting U.S. presidential elections: one is leveraging spatial information effectively to predict election outcomes with considering the impact of spatial factors, and the other is the limited data from swing states. Finding ways to use information from other states to analyze spatial factors and predict the final result is crucial to overcoming these challenges.
@@
Spatial dependence among spatial units exists widely in various fields including economics (Anselin,, 1988), finance (Grennan,, 2019), political science (Peters et al.,, 2004), ecology (Ver Hoef et al.,, 2018) and climate sciences (Okunlola et al.,, 2021) and so on. The spatial autoregressive (SAR) model (Cliff and Ord,, 1973; Ord,, 1975; Cliff and Ord,, 1975) and its variants are extensively applied in numerous empirical studies concerning spatial competition and spatial spillover effects. For instance, the SAR model is employed to investigate the crime rates in 49 areas of Columbus, Ohio (Anselin,, 1988), budget spillover effects and policy interdependence (Case et al.,, 1993), cigarette demands in various states in the U.S. (Baltagi and Li,, 2004), agricultural land prices in Northern Ireland (Kostov,, 2009), the mutual influence on the stock returns of companies with the same headquarters location (Pirinsky and Wang,, 2006) and the group effect of the dividend policy of listed companies (Grennan,, 2019).
@@
When the sample size of the target data is limited, it is possible to obtain less precise estimates for most statistical models. This issue is particularly relevant in spatial econometrics, where sample sizes can often be quite limited. For instance, in the study investigating crime rates in Columbus, Ohio, the sample size is restricted to just 49 (Anselin,, 1988). This limitation underscores the necessity for additional research aimed at leveraging information from similar datasets to elevate estimation performance. Consequently, there is a pressing need for the exploration of methods to improve the estimation performance of the SAR models through the incorporation of information derived from similar source datasets. Transfer learning is a useful approach to incorporate information from source datasets (Pan and Yang,, 2009; Torrey and Shavlik,, 2010; Zhuang et al.,, 2021; Hu and Zhang,, 2023). The main idea of transfer learning is that the knowledge acquired in one context can be transferred to improve the performance of estimation or prediction when applied to another, even if the two tasks or domains are not identical. In transfer learning, there are typically two main components: (1) The first one is the target domain, to which the knowledge is transferred. The target domain is of our interest but particularly its sample size is limited. (2) The second one is the source domain, from which the knowledge is transferred. In general, the model is pre-trained using the source datasets.
@@
In the recent literature, Bastani, (2021) introduced a two-step estimator that employs high-dimensional techniques to capture biases between the target and source data and has provided an upper bound for the proposed estimators. Li et al., 2022b considered transfer learning for estimation and prediction of a high dimensional linear regression and proposed a novel data-driven Trans-Lasso method for detecting the transferable sources. Tian and Feng, (2023) introduced a novel entropy-based method for identifying transferable sources and further constructed confidence intervals for each coefficient under the generalized linear model. For more reference on related works, see Cai and Pu, (2019), Cai et al., 2024a ; Cai et al., 2024b , and Li et al., 2022c ; Li et al., 2022a .
@@
In this paper, in order to deal with two aforementioned challenges, spatial dependence and sample limitation, in U.S. presidential election prediction problems, we present a transfer learning framework within the SAR model, which we refer to as tranSAR, designed to improve both estimation and prediction. Our study makes three contributions from methodology, statistical theory and empirical application in the literature of the SAR models. First, in methodology, we introduce a two-stage algorithm ($\mathcal{A}$-TranSAR), consisting of a transferring stage and a debiasing stage, to estimate the unknown parameters when the informative source data indexed by $\mathcal{A}$ are known. If we do not know which sources to transfer, a transferable source detection algorithm is proposed to detect informative sources data based on spatial residual bootstrap to retain the spatial dependence. To the best of our knowledge, this is the first work to introduce the transfer learning to the SAR models. Second, in statistical theory, we establish the theoretical convergence rates for the $\mathcal{A}$-TranSAR estimators under the SAR model and theoretically show that the transfer learning can improve the estimation accuracy with the help of the information from the transferable source data. We also derive the detection consistency of the transferable source detection algorithm to avoid the negative transfer when which sources to transfer is unknown. Third, in the empirical application, we apply our method to predict the election outcomes in swing states in the U.S. presidential election via utilizing polling data from the last U.S. presidential election along with other demographic and geographical data. The empirical results show that our method outperforms traditional spatial econometrics methods.
Here, we emphasize the differences between our work and Tian and Feng, (2023). Although our work shares a similar idea of transfer leaning with the previous literature, there are at least two significant differences from both theoretical and algorithmic perspectives. (1) Tian and Feng, (2023) is based on generalized linear models for independent data. In contrast, our research focuses on spatially dependent data, which brings significant challenges for theoretical guarantees. This is because previous theories heavily rely on concentration inequalities applicable to independent data, and few concentration inequalities have been established for spatially dependent data. To address this, we develope a new theoretical foundation tailored to spatially correlated data. (2) For the algorithmic aspect, the transferable detection method in Tian and Feng, (2023) relies on sample splitting to identify transferable sources, a technique that is inapplicable for spatial data as spatial data are not independent. To this end, we propose a new method based on spatial bootstrap and establish its theoretical guarantees.
In the rest of our paper, we firstly introduce the real application problem (predicting the U.S. presidential election result) and the associated transfer learning framework for the SAR model in Section 2. Then, we introduce its estimation and source detection algorithm in Section 3. In Section 4, we evaluate the finite sample performance of our proposed method via simulation studies. Further, the application to predict U.S. presidential elections with transfer learning is conducted in Section 5. Section 6 establishes the estimation convergence rates of the transfer learning estimators for the SAR model and derive the detection consistency of the transferable source detection algorithm. We conclude the paper in Section 7. Additional simulation results, theoretical analysis, and all proofs are provided in the supplementary materials.
@@
2 Problem and Model Settings
In this work, we focus on predicting the U.S. presidential election results of each county in swing states by incorporating the spatial relationship among counties. We try to predict the election results in 2020 and 2024 elections using the polls from the 2016 and 2020 election U.S. presidential election and some other demographic data and geographic data, respectively. It is worth noting that swing states have undergone historical shifts over time. The process of identifying swing states in previous elections typically involves an assessment of the closeness of the vote margins in each state. This analysis incorporates historical election outcomes, opinion polls, political trends, recent developments since the previous election, and the specific attributes, strengths, or vulnerabilities of the candidates in contention. In Section 2.1, we describe the motivation problem and data. Here, we reiterate the challenges we face, namely the utilization of spatial information and data scarcity. To this end, we propose a novel transfer learning framework based on the SAR model aimed at addressing above mentioned two challenges.
2.1 Problem and Data
Problem:
We aim to handle the challenges of utilizing spatial information and analyzing spatial factors in predicting the U.S. presidential election. From Figure 1, it is evident that the county-level support rates in the U.S. presidential election display a spatial relationship. Spatial relationships among the data are also prevalent, making it essential to incorporate spatial information into the analysis. In U.S. presidential elections, the distribution of votes exhibits significant spatial dependency, a phenomenon particularly obvious over counties. Specifically, the voting patterns of a county tend to align closely with those of its neighboring counties. This similarity can often be attributed to the shared socio-economic characteristics that arise from geographical proximity, such as industrial structure, income levels, and educational attainment (Fotheringham et al.,, 2021; Kim et al.,, 2003). Additionally, whether a county is urban or rural plays a substantial role in this spatial dependency. Urban counties generally show stronger support for the Democratic Party, whereas suburban and rural counties tend to favor the Republican Party. Thus, in analyzing voting distributions, the urban or rural status of a county becomes a critical factor, reflecting not only the geographical distribution of political preferences but also highlighting the political and social divides between urban and rural areas. Recognizing this spatial dependency helps in more accurately predicting election outcomes and in forming strategies for political parties. Thus the first challenge lies in how to incorporate spatial information into transfer learning techniques.
Figure 1: Results by county of the 2020 United States presidential election (Magog the Ogre,, 2020). The red areas represent support for Trump, while the blue areas indicate support for Biden. The intensity of the color corresponds to the degree of support.
Though the method proposed by Tian and Feng, (2023) provides a valuable framework to transfer learning and has been applied to the U.S. presidential election, their method is for independent data and does not account for spatial information during transfer learning. Furthermore, the second challenge is the issue of data scarcity. It is evident that the average sample size per state (the number of counties) is only around 60. Relying only on data from swing states for spatial analysis could harm the accuracy of the estimates.
Data sources and processing:
The digital boundary definitions of the United States congressional district are available at https://cdmaps.polisci.ucla.edu/, and the cartographic boundary shape files of counties in the USA are available at the United States Census Bureau (https://www2.census.gov/geo/tiger/). The spatial weight matrix is constructed using above two geographic datasets. Here we apply the “queen” contiguity-based spatial weights and row-normalized. The “queen” criterion defines neighbors as spatial units sharing a common edge or a common vertex. The response is the difference of support at the county level for presidential candidates from both parties in 2016 and 2020, available at https://github.com/tonmcg/US_County_Level_Election_Results_08-20. The county-level demographic information forms the predictors, collected at https://data.census.gov/. Among all states and the federal district, we exclude Alaska, Hawaii, and Washington, D.C.. In these 48 states, there are 3107 counties and we have 142 county-level predictors.
2.2 Model Settings
In order to deal with the two aforementioned challenges in U.S. presidential election prediction problems, we present a transfer learning framework within the spatial autoregressive (SAR) model. For a pre-specified swing state as the target data, we consider the following SAR model,
$Y^{(0)}=\sum_{l=1}^{p}\lambda_{l0}W_{l}^{(0)}Y^{(0)}+\sum_{j=1}^{q}X^{(0)}_{j}\beta_{j0}+V^{(0)},$ (1)
@@ 2.2 Model Settings
where the difference of support $Y^{(0)}$ is an $n_{0}\times 1$ vector of spatially correlated dependent variables, $W_{l}^{(0)}$ is an $n_{0}\times n_{0}$ known non-stochastic spatial weight matrix for each $l\in[p]$, the county-level demographic information $X_{j}^{(0)}$ is an $n_{0}\times 1$ vector of exogenous regressors independent of the error term $V^{(0)}$ for each $j\in[q]$, $V^{(0)}$ is an $n_{0}\times 1$ vector of independently and identically distributed (i.i.d.) random variables with mean 00 and variance $\sigma^{2}>0$. Let $\lambda_{0}={(\lambda_{10},\dots,\lambda_{p0})}^{\intercal}$, $\beta_{0}=(\beta_{10},\dots,\beta_{q0})^{\intercal}$ and $\theta_{0}\equiv$ ${({\lambda_{0}}^{\intercal},{\beta_{0}}^{\intercal})}^{\intercal}$.
@@
Similar to general transfer learning settings as in Bastani, (2021), Li et al., (2021), Tian and Feng, (2023), in addition to the target study (1), additional samples are obtained from $K$ auxiliary studies (also termed as source datasets) from other states, where $K$ is assumed to be fixed. For the $k$-th source dataset, we also consider the SAR model,
$Y^{(k)}=\sum_{l=1}^{p}\lambda_{l0}^{(k)}W_{l}^{(k)}Y^{(k)}+\sum_{j=1}^{q}X^{(k)}_{j}\beta^{(k)}_{j0}+V^{(k)},\ k\in[K],$ (2)
@@
$\mathcal{A}\equiv\mathcal{A}_{h}=\{0\leq k\leq K:\|\delta^{(k)}\|_{1}\leq h\},$ (3)
@@ 3.1 Estimation with known informative source data
$\widehat{\omega}=\underset{\omega\in\mathbb{R}^{p+q}}{\arg\min}\left\{\frac{1}{2n_{\mathcal{A}}}\sum_{k\in\mathcal{A}}\operatorname{\mathit{\ell}}_{1}\left(\mathfrak{D}^{(k)}\mid\omega\right)+\lambda_{\omega}\|\omega\|_{1}\right\},$ (4)
where $\operatorname{\mathit{\ell}}_{1}(\cdot\mid\omega)$ is a generic loss function with parameter $\omega$, and $n_{\mathcal{A}}\equiv\sum_{k\in\mathcal{A}}n_{k}$ is the total sample size of all informative source data. For the SAR model, the estimator could be the two-stage least squares (2SLS) estimator (Kelejian and Prucha,, 1998; Lee,, 2003), generalized method of moments estimator (GMM) (Kelejian and Prucha,, 1999; Lee,, 2007; Liu and Lee,, 2010), quasi-maximum likelihood estimator (Lee,, 2004, QMLE) and so on. The main requirement for the first step estimator is a suitable convergence rate discussed in detail in Section 6. Here the tuning parameter $\lambda_{\omega}$ could be zero in low dimensional cases. When the dimension of parameters is greater than the sample size in high dimensional cases, the penalized SAR estimation with nonzero $\lambda_{\omega}$ could produce a sparse estimator(Higgins and Martellosio,, 2023). See also Zhang and Yu, (2018) for spatial weight matrices selection with model averaging, and Tibshirani, (1996); Fan and Li, (2001); Zhang, (2010); Wang et al., (2011); Su et al., (2017) for more detail in regularization method.
@@
The index set $\mathcal{A}$ may or may not contain the index 00. This depends on the actual application scenarios. In some cases the source data are available, and in the other cases, it may be difficult to obtain due to data privacy protection and security considerations. In the latter case, the owner of the source data can provide the estimation results obtained by the first stage learner without the target data instead of the original source data, and the target task will use this result for the next debiasing stage.
@@
$\underset{\delta\in\mathbb{R}^{p+q}}{\arg\min}\frac{1}{2n_{0}}\operatorname{\mathit{\ell}}_{2}\left(\mathfrak{D}^{(0)}\mid\widehat{\omega}+\delta\right)+\lambda_{\delta}\|\delta\|_{1}.$ (5)
@@
$\widehat{\theta}_{\mathcal{A}\textit{-TranSAR}}=\widehat{\omega}+\widehat{\delta}.$ (6)
@@
Input: target dataset: $\mathfrak{D}^{(0)}$; sources datasets: $(\mathfrak{D}^{(k)},k\in\mathcal{A})$; transferring set: $\mathcal{A}$; and tuning parameters: $\lambda_{\omega}$ and $\lambda_{\delta}$